# Cramming 1568 Tokens into a Single Vector and Back Again: Exploring the Limits of Embedding Space Capacity

### v1

@@ Abstract
A range of recent works addresses the problem of compression of sequence of tokens into a shorter sequence of real-valued vectors to be used as inputs instead of token embeddings or key-value cache. These approaches allow to reduce the amount of compute in existing language models. Despite relying on powerful models as encoders, the maximum attainable lossless compression ratio is typically not higher than x10. This fact is highly intriguing because, in theory, the maximum information capacity of large real-valued vectors is far beyond the presented rates even for 16-bit precision and a modest vector size. In this work, we explore the limits of compression by replacing the encoder with a per-sample optimization procedure. We show that vectors with compression ratios up to x1500 exist, which highlights two orders of magnitude gap between existing and practically attainable solutions. Furthermore, we empirically show that the compression limits are determined not by the length of the input but by the amount of uncertainty to be reduced, namely, the cross-entropy loss on this sequence without any conditioning. The obtained limits highlight the substantial gap between the theoretical capacity of input embeddings and their practical utilization, suggesting significant room for optimization in model design.
@@ 1 Introduction
In the context of LLMs, the challenge of encoding prompts and long contexts is particularly important because of the quadratic computational cost of the self-attention mechanism in Transformers. Several works have explored the possibility of replacing token-based prompts with a smaller set of dense vectors (Lester et al., 2021; Li and Liang, 2021; Gao, 2024; Li et al., 2024), thereby shortening the input sequence and reducing the computational budget. These methods have demonstrated token-to-vector lossy compression ratios on the order of x500 with 8B-parameter models, indicating that it is possible to retain the critical information in a significantly reduced number of vectors. However, lossless compression is still limited by approximately factor of 10.
Furthermore, another important aspect of embeddings is that they also constitute the operational space in which the model’s computations take place through residual stream. In memory-augmented architectures (Weston et al., 2015; Sukhbaatar et al., 2015; Burtsev et al., 2021), these embeddings can act as additional storage or as a recurrent state for passing information between time steps (Dai et al., 2019; Bulatov et al., 2022; Chevalier et al., 2023; Behrouz et al., 2024), essentially serving as an episodic memory. Moreover, recent approaches have explored the power of latent space reasoning (Hao et al., 2024) where high-capacity embeddings enable models to perform complex multi-step tasks directly in latent space. Consequently, the capacity of these vectors is crucial not only for efficient input representation, but also for increasing the overall expressiveness and computational power of models (Merrill and Sabharwal, 2023; Strobl et al., 2024; Sanford et al., 2024). Better understanding of the latent capacity of input vectors, could significantly help to improve encoding and retrieving of contextual information, episodic memory, as well as complex reasoning within large language models.
@@
3. We show that the capacity limits remain consistent across different text lengths and domains, including natural text and randomly generated word sequences.
@@ 2 Related Work
The approach of compressing the context of LLMs into a shorter sequence of input or KV-cache vectors is actively explored for various purposes, yet no standardized terminology or unified methodology has emerged.
@@
One application for input compression is connected with efficient processing of long contexts with LLMs. RMT Bulatov et al. (2022) and AutoCompressors Chevalier et al. (2023) train the whole language model in a recurrent manner to compress the information from input segments to summary vectors and later reuse them to solve long-context tasks. Ge et al. (2023) use an autoencoder architecture with a frozen LLM as a decoder and adapt the same LLM for the encoder using LoRA Hu et al. (2022). The resulting pipeline is pretrained using autoencoding and language modeling objectives, and then finetuned for language tasks, achieving the effective compression rate of x4. SelfCP Gao (2024) uses the base LLM itself as a compressor using a trainable adapter to aggregate compressed states across multiple segments. 500xCompressor Li et al. (2024) extends the autoencoding approach with layer-wise connections and additional language pretraining tasks, exploring compression ratios up to x480, though at the cost of substantial quality degradation. In contrast, our method, applied to models of comparable size (up to 8 billion parameters), demonstrates that a compression rate of x1568 can be achieved with no loss in reconstruction quality
@@
Another line of work targets compressing only the prompts to reduce inference costs. Gist tokens Mu et al. (2023) are prompt representations compressed by the LLM itself, finetuned with a special mask. Gisting allows to achieve prompt compression rate up to x26 with only minor loss in model performance. LLMLingua Jiang et al. (2023a) decouples the compression operation from the LLM and introduces a coarse-to-fine prompt compression strategy with a budget controller and token-level iterative compression, achieving up to 20 times compression with negligible performance loss. Subsequent works Jiang et al. (2023b); Pan et al. (2024) extend this framework to long contexts, improving information retention through data distillation. Additionally, Morris et al. (2023) suggest that some information about prompts can be recovered from the model predictions themselves. In the current work we apply a per-sample optimization process instead to explore the fundamental limits of compression and establish upper bounds on compression rates that far exceed prior work.
@@
Language models have also been investigated for lossless text compression. LLMZip Valmeekam et al. (2023) improves standard compression by ranking candidates via next-token probabilities, while FineZip Mittu et al. (2024) accelerates compression through finetuning and dynamic context management for better efficiency. The theoretical capabilities of LLMs in compression pipelines can be estimated via measuring the bits-per-token metric over a representative textual corpus. Huang et al. (2024); Guo et al. (2024) provide such measurements for public LLMs and establish the connection between compression rate and model performance, measured by diverse benchmark scores. Unlike these methods, we do not rely on external compression algorithms. Instead, we achieve lossless compression using only the LLM itself, providing both theoretical insights and practical demonstrations of compression limits.
@@
Some works utilize the trainable input tokens approach in other ways. Burtsev et al. (2021) uses memory tokens as additional representation storage, Beltagy et al. (2020) and Zaheer et al. (2020) use similar global tokens to enhance long-range information flow. Li and Liang (2021); Lester et al. (2021); Gao (2024); Liu et al. (2022) explore trainable soft prompts for one or multiple layers as an alternative to finetuning model weights. Our findings about the representation capacity can represent the potential efficiency limits of such methods, based on how far the model behavior can be changed using trainable tokens.
@@ 3 Method
Trainable [mem] vectors are inspired by Memory Transformers Burtsev et al. (2021), but here these vectors are designed to encode an entire text sequence. The training method is borrowed from Prompt Tuning Lester et al. (2021), with only a set of special input embeddings optimized while all parameters of the language model are frozen.
@@
Let estimate an upper bound on the number of tokens that can be generated from a single input vector by a language model. The input vector has a dimension $d_{model}$, with each element represented by $b$ bits, so that the total information content is approximately $d_{model}\times b$ bits. Given a vocabulary of size $|\mathcal{V}|$, where each token carries at most $\log_{2}|\mathcal{V}|$ bits of information, the maximum number of tokens $L$ that can be generated is bounded by:
@@ 4 Experiments and Results
We use models from Pythia suite (160M, 410M, 1.4B, 2.8B) Biderman et al. (2023), OPT-1.3B Zhang et al. (2022), OLMo-1B Groeneveld et al. (2024), Sheared-LLaMA-1.3B Xia et al. (2024), and Llama-3 models (1B, 3B, 8B) Dubey et al. (2024). List of all used models with links to HuggingFace Hub are in Appendix A.
@@
4.1 Decoding Capacity of a Single Token
@@ 4.3 Sensitivity of Compression to Text Complexity
For 50 text samples from the PG-19 at each target length (ranging from 8 up to 1568 tokens, and to 3072 for larger models) we measured cross-entropy both before ($H_{\text{LM}}$) and after ($H_{\text{LM+[mem]}}$) compression (see Eq. 4). Figure 3 compares results across multiple models, and full results are provided in Appendix C.
On Fig. 3 the models demonstrate linear relationship between cross-entropy before and after compression for not perfectly compressible texts (i.e., lying above the red dotted line), indicating constant value of information gain (or, reduction in cross-entropy). Texts with cross-entropy smaller than a model’s information gain are perfectly reconstructed.
@@ 4.5 Embedding Capacity Utilization
In Fig. 5 (top), when comparing all models with roughly 1B parameters, there are two groups: older models (e.g., OPT and Pythia) show lower capacity utilization, whereas newer models (e.g., Llama, ShearedLlama, and OLMo) demonstrate higher utilization despite having the same theoretical capacity. This disparity indicates that the quality of pre-training (data, compute budget) influences the extent to which a model can exploit its input vectors capacity.
@@
In Fig. 5 (bottom), the Pythia models show an interesting trend: as model size increases, capacity utilization decreases. This pattern suggests that the larger Pythia models may be under-trained relative to their theoretical potential. In contrast, Llama and OLMo models show higher capacity utilization. Based on these observations, we hypothesize that capacity utilization could serve as an indicator of the pre-training status and guide further training.
@@ 5 Discussion and Conclusions
While our experiments push the boundaries of compression with LLMs and offer insights into their upper capacity limits, the nature of the obtained representations remains largely unclear. We have analyzed the structure of the space of trained [mem] vectors in Appendix D, but more in-depth analysis is needed to determine the semantic properties of the vectors and their potential value in downstream tasks. Our findings are limited to Transformer-based models with up to 8 billion parameters due to computational constraints. Investigating the representation space of larger models, as well as exploring alternative architectures such as recurrent and memory-augmented models, remains an important avenue for future research. In our study with different text sources, we generate random text by sampling words from a dictionary. While this approach simplifies the analysis, it may slightly overestimate model capacity compared to sampling directly from a tokenizer’s vocabulary, as dictionary words can be split on multiple tokens by model.
@@ Appendix B Collecting Texts from the Fanfics Library
Appendix C Results of Evaluating Text Compression for All Models
Here we provide results for all evaluated models in Fig. 6. Results are discussed in Section 4.3.
Appendix D Understanding the Structure of Compressed Vectors

### v2

@@ Abstract
A range of recent works addresses the problem of compression of sequence of tokens into a shorter sequence of real-valued vectors to be used as inputs instead of token embeddings or key-value cache. These approaches are focused on reduction of the amount of compute in existing language models rather than minimization of number of bits needed to store text. Despite relying on powerful models as encoders, the maximum attainable lossless compression ratio is typically not higher than x10. This fact is highly intriguing because, in theory, the maximum information capacity of large real-valued vectors is far beyond the presented rates even for 16-bit precision and a modest vector size. In this work, we explore the limits of compression by replacing the encoder with a per-sample optimization procedure. We show that vectors with compression ratios up to x1500 exist, which highlights two orders of magnitude gap between existing and practically attainable solutions. Furthermore, we empirically show that the compression limits are determined not by the length of the input but by the amount of uncertainty to be reduced, namely, the cross-entropy loss on this sequence without any conditioning. The obtained limits highlight the substantial gap between the theoretical capacity of input embeddings and their practical utilization, suggesting significant room for optimization in model design.
@@ 1 Introduction
In the context of LLMs, the challenge of encoding prompts and long contexts is particularly important because of the quadratic computational cost of the self-attention mechanism in Transformers. Several works have explored the possibility of replacing token-based prompts with a smaller set of dense vectors (Lester et al., 2021; Li and Liang, 2021; Gao, 2024; Li et al., 2024b), thereby shortening the input sequence and reducing the computational budget. These methods have demonstrated token-to-vector lossy compression ratios on the order of x500 with 8B-parameter models, indicating that it is possible to retain the critical information in a significantly reduced number of vectors. However, lossless compression is still limited by approximately factor of 10.
In memory-augmented architectures (Weston et al., 2015; Sukhbaatar et al., 2015; Burtsev et al., 2021), these embeddings act as additional storage or as a recurrent state for passing information between time steps (Dai et al., 2019; Bulatov et al., 2022; Chevalier et al., 2023; Behrouz et al., 2024), essentially serving as an episodic memory. Moreover, recent approaches have explored the power of latent space reasoning (Hao et al., 2024) where high-capacity embeddings enable models to perform complex multi-step tasks directly in latent space. Consequently, the capacity of these vectors is crucial not only for efficient input representation, but also for increasing the overall expressiveness and computational power of models (Merrill and Sabharwal, 2023; Strobl et al., 2024; Sanford et al., 2024). Better understanding of the latent capacity of input vectors, could significantly help to improve encoding and retrieving of contextual information, episodic memory, as well as complex reasoning within large language models.
@@
3. We show that the capacity limits remain consistent across different text lengths and domains, including natural text and random word sequences.
@@ 2 Related Work
Approaches to compressing LLM context into a shorter sequence of input or KV-cache vectors are explored for various purposes, yet no standardized terminology or unified methodology has emerged.
@@
One application for input compression is connected with efficient processing of long contexts with LLMs. RMT Bulatov et al. (2022) and AutoCompressors Chevalier et al. (2023) train the whole language model in a recurrent manner to compress the information from input segments to summary vectors and later reuse them to solve long-context tasks. ICAE Ge et al. (2024) uses an autoencoder architecture with a frozen LLM as a decoder and adapt the same LLM for the encoder using LoRA Hu et al. (2022). The resulting pipeline is pretrained using autoencoding and language modeling objectives, and then finetuned for language tasks, achieving the effective compression rate of x4. SelfCP Gao (2024) uses the base LLM itself as a compressor using a trainable adapter to aggregate compressed states across multiple segments. 500xCompressor Li et al. (2024b) extends the autoencoding approach with layer-wise connections and additional language pretraining tasks, exploring compression ratios up to x480, though at the cost of substantial quality degradation. Alternative approaches aim to compress KV activations instead of input tokens. Some methods achieve this by estimating token relevance, either through training-free Zhang et al. (2023); Li et al. (2024a) or training-based approaches Qin and Van Durme (2023); Qin et al. (2024), to prune irrelevant tokens and focus computation on the most informative ones. These strategies can yield high-quality but lossy compression with ratios up to x20. This result can be improved by finetuning models to leverage the resulting cache representations more effectively. This way, KV-Distill Chari et al. (2025) can reduce cache size up to 100 times with negligible loss in QA performance. In contrast, our method, applied to models of comparable size (up to 8 billion parameters), demonstrates that a compression rate of x1568 can be achieved without any loss in reconstruction quality.
@@
Another line of work targets compressing only the prompts to reduce inference costs. Gist tokens Mu et al. (2023) are prompt representations compressed by the LLM itself, finetuned with a special mask. Gisting allows to achieve prompt compression rate up to x26 with only minor loss in model performance. LLMLingua Jiang et al. (2023) decouples the compression operation from the LLM and introduces a coarse-to-fine prompt compression strategy with a budget controller and token-level iterative compression, achieving up to x20 compression with negligible performance loss. Jiang et al. (2024); Pan et al. (2024) extend this framework to long contexts, improving information retention through data distillation. Additionally, Morris et al. (2024) suggest that some information about prompts can be recovered from the model predictions themselves. In the current work we apply a per-sample optimization process instead to explore the fundamental limits of compression and establish upper bounds on compression rates that far exceed prior work.
@@
Language models have also been investigated for lossless text compression. [^1] Delétang et al. (2024) explore data compression (texts to sequences of bits) using Arithmetic Coding and LLMs. This line of research is orthogonal to ours, as the sequence of bits cannot be passed as an input to an LLM. We analyze our approach from the data compression perspective in Appendix F. LLMZip Valmeekam et al. (2023) improves standard compression by ranking candidates using next-token probabilities, while FineZip Mittu et al. (2024) accelerates compression through fine-tuning and dynamic context management for better efficiency. The capabilities of LLMs in compression pipelines can be measured in bits-per-token over a representative textual corpus. Huang et al. (2024); Guo et al. (2024) provide such measurements for public LLMs and establish the connection between compression rate and model performance, measured by diverse benchmark scores. Unlike these methods, we do not rely on external compression algorithms. Instead, we achieve lossless compression using only the LLM itself, providing both theoretical insights and practical demonstrations of compression limits.
@@
Some works use the trainable input tokens in other ways. Burtsev et al. (2021) uses memory tokens as additional representation storage, Beltagy et al. (2020); Zaheer et al. (2020) use similar global tokens to enhance long-range information flow. Li and Liang (2021); Lester et al. (2021); Gao (2024); Liu et al. (2022) explore trainable soft prompts as an alternative to finetuning model weights. Our findings about the representation capacity can represent the potential efficiency limits of such methods, based on how far the model behavior can be changed using trainable tokens.
@@ 3 Method
Trainable [mem] vectors are inspired by Memory Transformers Burtsev et al. (2021), but here these vectors are designed to encode an entire text sequence. The training method is similar to Prompt Tuning Lester et al. (2021), with only a set of special input embeddings optimized while all parameters of the language model are frozen.
@@
Let’s estimate an upper bound on the number of tokens that can be generated from a single input vector by a language model. The input vector has a dimension $d_{model}$, with each element represented by $b$ bits, so that the total information content is approximately $d_{model}\times b$ bits. Given a vocabulary of size $|\mathcal{V}|$, where each token carries at most $\log_{2}|\mathcal{V}|$ bits of information, the maximum number of tokens $L$ that can be generated is bounded by:
@@ 4 Experiments and Results
We use models from Pythia suite (160M, 410M, 1.4B, 2.8B) Biderman et al. (2023), OPT-1.3B Zhang et al. (2022), OLMo-1B Groeneveld et al. (2024), Sheared-LLaMA-1.3B Xia et al. (2024), Llama-3 models (1B, 3B, 8B) Dubey et al. (2024), and Mamba (130M, 370M, 790M, 1.4B) Gu and Dao (2024). List of all used models with links to HuggingFace Hub are in Appendix A.
@@
4.1 Decoding Capacity of a Single Vector
@@ 4.3 Sensitivity of Compression to Text Complexity
For 50 text samples from the PG-19 at each target length (ranging from 8 up to 1568 tokens, and to 3072 for larger models) we measured cross-entropy both before ($H_{\text{LM}}$) and after ($H_{\text{LM+[mem]}}$) compression (see Eq. 4). Figure 3 compares results across Pythia and LLama models, and full results for all models are provided in Appendix D.
In Fig. 3, the models demonstrate linear relationship between cross-entropy before and after compression for not perfectly compressible texts (i.e., lying above the red dotted line), indicating constant value of information gain (or, reduction in cross-entropy). Texts with cross-entropy lower than a model’s information gain are perfectly reconstructed.
@@ 4.5 Embedding Capacity Utilization
In Fig. 5 (top), when comparing all models with roughly 1B parameters, there are two groups: (1) older models (e.g., OPT and Pythia) show lower capacity utilization, whereas (2) newer models (e.g., Llama, ShearedLlama, Mamba, and OLMo) demonstrate higher utilization despite having the same theoretical capacity. This disparity indicates that the quality of pre-training (data, compute budget, improvements in architecture) influences the extent to which a model can exploit its input vectors capacity.
@@
In Fig. 5 (bottom) we can see three groups: (1) and (2) same as on top plot with older and newer models, and (3) group of Mamba models. The Pythia models show an interesting trend: as model size increases, capacity utilization decreases. This pattern suggests that the larger Pythia models may be under-trained relative to their theoretical potential. In contrast, Llama and OLMo models show higher capacity utilization. Based on these observations, we hypothesize that capacity utilization could serve as an indicator of the pre-training status and guide further training.
4.6 Non-Transformer Models: Mamba
Notably, the findings from the previous sections are not limited to Transformer-based language models. State-space Mamba models can also encode an entire text into a single [mem] vector and perfectly reconstruct the text when its cross-entropy falls below the model’s capacity threshold, e.g., Mamba-1.4B can encode texts of $512$ tokens into a single [mem] vector, similar to other 1B models (LLama-3.2-1B, Sheared-LLaMa-1.3B). This demonstrates that obtained results are architecture-agnostic (more details on Mamba results in Appendices C, 6 and 3).
On capacity utilization (Fig. 5), the Mamba models exhibit a consistent pattern across both panels that depends on model size. In the top plot, the increase in correctly decoded tokens is nearly linear with respect to each model’s theoretical capacity. This implies that each additional parameter in the embedding space translates into greater decoding accuracy at a nearly constant rate. In the bottom plot, capacity utilization rises monotonically from the 130M to the 1.4B model. This is the opposite of the downward trend seen in Pythia and shows that larger Mambas progressively make better use of their learned input space.
@@ 5 Discussion and Conclusions
While our experiments push the boundaries of compression with LLMs and offer insights into their upper capacity limits, the nature of the obtained representations remains largely unclear. We have analyzed the structure of the space of trained [mem] vectors in Appendix E, but more in-depth analysis is needed to determine the semantic properties of the vectors and their potential value in downstream tasks. Our findings are limited to Transformer-based and Mamba models with up to 8 billion parameters due to computational constraints. Investigating the representation space of larger models, as well as exploring alternative architectures such as recurrent and memory-augmented models, remains an important avenue for future research. In our study with different text sources, we generate random text by sampling words from a dictionary. While this approach simplifies the analysis, it may slightly overestimate model capacity compared to sampling directly from a tokenizer’s vocabulary, as dictionary words can be split on multiple tokens by model.
@@
Acknowledgments
We are thankful to SberDevices for granting us access to additional computational resources. This work was partially supported by the Ministry of Economic Development of the Russian Federation (code 25-139-66879-1-0003).
@@ Appendix A Models and Training Details
Mamba-130M state-spaces/mamba-130m-hf 0.13 768 50280
Mamba-370M state-spaces/mamba-370m-hf 0.38 1024 50280
Mamba-790M state-spaces/mamba-790m-hf 0.79 1536 50280
Mamba-1.4B state-spaces/mamba-1.4b-hf 1.4 2048 50280
@@ Appendix B Collecting Texts from the Fanfics Library
Appendix C Non-Transformer Language Models: Mamba
Mamba (Gu and Dao, 2024) is a state space language model (SSM) (Gu et al., 2022), in contrast to attention-based Transformers. We applied the same compression procedure used for the Transformer models and evaluated the official Mamba checkpoints from the Hugging Face Hub (Table 2). We found that Mamba can also encode texts into a single [mem] vector and successfully reconstruct them whenever the text’s cross-entropy falls below model’s capacity threshold, indicating the effect is not unique to only Transformer models (Figs. 6 and 3).
Appendix D Results of Evaluating Text Compression for All Models
Here we provide results for all evaluated models in Figs. 6 and 3. Results are discussed in Section 4.3.
Figure 6: Information gain of text compression to [mem] vector doesn’t depend on language understanding capabilities of models. Compression results for various language models show the relationship between the cross-entropy (CE) of the original and decompressed texts. If the text CE falls below a model-specific threshold (red line), the text is losslessly compressed. This value is a input vector capacity in terms of entropy (Information Gain, $C_{H}$). For texts that are not perfectly compressed, the compression process reduces their CE to a consistent, model-specific value (bias of the black dashed line). Larger models (e.g., Llama-3.1-8B) can handle longer texts before reaching the compression threshold, due to their greater capacity compared to smaller models (e.g., Pythia-160M). This behavior holds for both natural texts (PG-19) and unnatural random texts consisting of random word sequences. The result is not limited to transformer-based architectures: non-transformer models, such as Mamba, exhibit the same pattern.
Model PG-19 Fanfics Random
Max, tok. Gain, tok. Information Gain Max, tok. Gain, tok. Information Gain Max, tok. Gain, tok. Information Gain
Pythia-160M 80 70.9$\pm$11.0 396.4$\pm$46.0 80 70.9$\pm$10.5 378.1$\pm$45.9 65 61.3$\pm$6.6 500.8$\pm$38.9
Pythia-410M 96 81.3$\pm$12.0 431.4$\pm$51.6 96 81.2$\pm$11.6 429.8$\pm$46.2 72 76.9$\pm$8.7 630.4$\pm$65.2
Pythia-1.4B 160 158.0$\pm$29.1 792.8$\pm$143.4 192 152.9$\pm$28.0 776.9$\pm$132.5 139 144.4$\pm$17.5 1108.2$\pm$136.2
Pythia-2.8B 128 150.1$\pm$50.7 740.3$\pm$234.9 - - - 141 134.5$\pm$24.7 1026.3$\pm$211.4
OPT-1.3B 128 132.2$\pm$23.8 712.8$\pm$143.3 - - - 116 129.3$\pm$16.4 1068.0$\pm$181.3
OLMo-1B 384 406.3$\pm$61.7 1901.0$\pm$254.5 - - - 249 257.3$\pm$55.0 1852.2$\pm$395.0
Sh.LLaMa-1.3B 512 383.6$\pm$38.4 1835.1$\pm$162.9 - - - 382 315.1$\pm$35.1 1893.0$\pm$210.0
Llama-3.2-1B 512 426.2$\pm$79.2 2119.9$\pm$364.8 512 449.6$\pm$83.7 2213.8$\pm$365.8 316 294.9$\pm$64.8 2265.2$\pm$498.7
Llama-3.2-3B 1024 720.3$\pm$80.2 3292.2$\pm$320.0 1024 734.1$\pm$85.0 3354.5$\pm$344.9 460 456.9$\pm$72.1 3382.6$\pm$585.2
Llama-3.1-8B 1568 1094.1$\pm$127.6 4865.7$\pm$546.6 1568 1071.8$\pm$168.6 4768.9$\pm$622.6 792 623.2$\pm$97.3 4541.2$\pm$758.6
Mamba-130M 32 65.5$\pm$18.1 371.4$\pm$94.6 - - - 84 62.6$\pm$17.1 535.7$\pm$121.8
Mamba-370M 128 113.0$\pm$27.6 585.3$\pm$131.1 - - - 106 94.5$\pm$25.5 743.2$\pm$186.9
Mamba-790M 256 204.2$\pm$33.0 1011.6$\pm$138.8 - - - 198 187.0$\pm$26.3 1421.3$\pm$179.1
Mamba-1.4B 512 348.9$\pm$42.9 1599.5$\pm$164.5 - - - 288 282.7$\pm$36.5 2062.3$\pm$257.3
Table 3: Compression capacity across different text sources and for all evaluated models. We report Decoding Capacity (in Tokens) ("Max, tokens" in the Table), Token Gain, and Information Gain for texts from PG-19, fanfics, random.
Appendix E Understanding the Structure of Compressed Vectors